∫dx/(e x + 4e -x ) = f (x) + c then f (x) is equal to

∫dx/(ex + 4e-x) = f (x) + c then f (x) is equal to

Grade:12

1 Answers

Gurwinder Kaur
12 years ago

∫dx/(ex + 4e-x)

=∫exdx/(ex2 + 4)  

=1/2∫2exdx/(ex2 + 4)

=1/2 log [ex2 + 4]+c

therefore,

fx= 1/2 log [ex2 + 4]

plz approve if u like..